Find the area of the largest rectangle with sides parallel to the coordinate axes which can be inscribed in the bounded by the two parabolas y=26-x^2 and y=x^2+2


1
Expert's answer
2021-05-17T13:04:02-0400

The graph of the above function is-




Each side intersect both functions twice


so area of rectangle = length * width


length = "(26-x^2) - (x^2+2) = 24 - 2x^2"

width = "-2x"


area of rectangle "= (24-2x^2)\times -2x"


"A=4x^3-48x"



taking the first derivative of the area


"A'=12x^2 - 48"


Putting "A'=0, 12x^2=48\Rightarrow x^2=4\Rightarrow x=\pm 2"


Again differentiating A'=


"A''=24x
\\

\text{
at } x=-2, A''=-48<0"


So Area is maximum at x=-2


Maximum area ="4(-2)^3-48(-2)=-32+96=64"